OpenStudy (1645323):

what is the decode of this TeX?TeX source: \frac{3}{x^{2}-4}

OpenStudy (phi):

put it between \ [ (with no space between those symbols) and \ ] \[ \frac{3}{x^{2}-4} \]

OpenStudy (1645323):

ok thanks

OpenStudy (1645323):

can you decode this one also? \{\frac{1}{3}, \frac{1}{2} , \frac{1}{3} , \frac{1}{4} \}

OpenStudy (1645323):

is it (1/3),1/2,1/3,1/4?

OpenStudy (1645323):

@phi

OpenStudy (phi):

yes. You used \{ do the same thing but replace the { with [